Please help me solve this problem.
Three times the measure of a complement of a certain angle is equal to 30 more than the measure of the supplement of that angle. Find the measure of each angle.

Answer by malakumar_kos@yahoo.com(315)


Three times the measure of a complement of a certain angle is equal to 30 more than the measure of the supplement of that angle. Find the measure of each angle.
Let the angle be = x
its complement = 90-x
its supplement = 180-x
three times the complement = 30 more than its supplement
3(90-x) = 30+(180-x)
270 - 3x = 30+180-x
270-180-30 = 3x-x
270-210 = 2x
60 = 2x
x = 60/2 = 30
The required two angles are 1)90-x = 90-30 = 60
2)180-x = 180-30 = 150